OpenStudy (unklerhaukus):

Solve the ODE \[(3x-y+1)\text d x-(6x-2y-3)\text d y=0\]

OpenStudy (unklerhaukus):

\[(3x-y+1)\text d x-(6x-2y-3)\text d y=0\] \[(3x-y+1)\text d x=(6x-2y-3)\text d y\] \[\frac{\text d y}{\text d x}=\frac{3x-y+1}{6x-2y-3}\] \[2\frac{\text d y}{\text d x}=\frac{6x-2y+2}{6x-2y-3}\] \[\text {let } w=6x-2y\] \[\frac{\text d w }{\text d x}=6-2\frac{\text d y }{\text dx}\] \[2\frac{\text d y}{\text d x}=6-\frac{\text d w }{\text d x}\] \[2\frac{\text d y}{\text d x}=\frac{w+2}{w-3}\] \[6-\frac{\text d w }{\text d x}=\frac{w+2}{w-3}\] \[6-\frac {w+2} {w-3}=\frac{\text d w }{\text d x}\] \[\frac {6(w-3)-w+2} {w-3}=\frac{\text d w }{\text d x}\] \[\frac {5w-16} {w-3}=\frac{\text d w }{\text d x}\] \[\text d x=\frac{w-3}{5w-16}\text d w\] \[\int\text d x=\int\frac{w-3}{5w-16}\text d w\]

OpenStudy (unklerhaukus):

that is all i got so far.

OpenStudy (unklerhaukus):

\[\ln \text dx =\int\frac{w-3}{5w-16}\text d x\]

OpenStudy (anonymous):

Making u = 3x - y maybe easier. If I did it correctly, you should end up with: \[\frac{du}{dx} = \frac{5u-2}{2u-3}\]So the integral becomes: \[ \int \frac{2u -3}{u-2}du = \int 5dx \implies \int (\frac{1}{u-2} + 2)du = 5x + C\]As for your integral, I think both doing long division and partial fractions should solve it.

OpenStudy (unklerhaukus):

those are the worst i will re do this with u = 3x - y

OpenStudy (anonymous):

I agree. I always disliked partial fractions. But it seems that they happen all the time when we are solving DE :-(

OpenStudy (unklerhaukus):

takes up to much paper

OpenStudy (unklerhaukus):

\[\text {let } w= 3x - y\]\[\frac{\text d w }{\text d x}=3-\frac{\text d y }{\text dx}\]\[\frac{\text d y}{\text d x}=3-\frac{\text d w }{\text d x}\]\[3-\frac{\text d w }{\text d x}=\frac{w+1}{2w-3}\]\[\frac{\text d w }{\text d x}=3-\frac{w+1}{2w-3}\]\[\frac{\text d w }{\text d x}=\frac{3(2w-3)- w+1}{2w-3}\]\[\frac{\text d w }{\text d x}=\frac{5w-8}{2w-3}\]

OpenStudy (anonymous):

Stupid me, it should be \(\large \frac{5(u-2)}{2u - 3} \). And, I got: \[\ u + 1 - (-\frac{du}{dx} + 3)(6x - 3 - 2(3x - u)) = 0 \]After substituting in.

OpenStudy (anonymous):